a number consists of two digits.Tens digit is 3 more than the ones digit and number is 7 times the sum of its dihits.Find the number.
Answer by lwsshak3(11628)

a number consists of two digits.Tens digit is 3 more than the ones digit and number is 7 times the sum of its digits.Find the number.
**
let u=units digit
let t= tens digit
number=10t+u
t=u+3 (tens digit 3 more than ones digit)
t+u= sum of digits
..
10t+u=7(t+u)
10(u+3)+u=7(u+3)+7u
10u+30+u=7u+21+7u
3u=9
u=3
t=u+3=6
number: 63